# A hydrodynamic approach to reproduce multiple spinning vortices in horizontally rotating three-dimensional liquid helium-4

### v1

@@ Abstract
This paper reports a three-dimensional (3D) simulation of a rotating liquid, helium-4, using a two-fluid model with spin-angular momentum conservation. Our model was derived from the particle approximation of an inviscid fluid with residual viscosity. Despite the fully classical mechanical picture, the resulting system equations were consistent with those of the conventional two-fluid model. We consider bulk liquid helium-4 to be an inviscid fluid, assuming that the viscous fluid component remains at finite temperatures. As the temperature decreased, the amount of the viscous fluid component decreased, ultimately becoming a fully inviscid fluid at absolute zero. Weak compressibility is assumed to express the volume change because some helium atoms do not render fluid owing to BECs or change states because of local thermal excitation. One can solve the governing equations for an incompressible fluid using explicit SPH (smoothed particle hydrodynamics), simultaneously reproducing density fluctuations and describing the fluid in a many-particle system. We assume the following fluid-particle duality: a hydrodynamic interfacial tension between the inviscid and viscous components or a local interaction force between two types of fluid particles. The former can be induced in the horizontal direction when non-negligible non-uniformity of the particles occurs during forced two-dimensional rotation, and the latter is non-negligible when the former is negligible. We performed a large-scale simulation of 3D liquid helium forced to rotate horizontally using 32 GPUs. Compared with the low-resolution calculation using 2.4 million particles, the high-resolution calculation using 19.6 million particles showed spinning vortices close to those of the theoretical solution. We obtained a promising venue to establish a practical simulation method for bulk liquid helium-4.
@@ 1 Introduction
Considering this hypothesis, the author recently developed a two-fluid model with spin-angular momentum conservation [6, 7, 8]. The following overview provides a general introduction to this subject. The Navier-Stokes equation with conservation of rotational angular momentum around the axes of fluid particles was originally derived from the constitutive law of continuum mechanics by D. Condiff in 1964 [9]. Half a century later, K. Müller [10] developed a particle approximation of the Navier-Stokes equation with spin-angular momentum conservation using smoothed particle hydrodynamics (SPH) [11, 12] for mesoscale thermal flow. Considering these studies, the author adopted the Navier-Stokes equations with spin-angular momentum conservation for normal fluids and employed SPH to discretize the equation systems using particles. In other words, we describe liquid helium-4 as a system of numerous particle interactions composed of two distinct types of classical fluid particles that maintain their rotational angular velocity around their axes under the assumption of rigid body rotations; this is referred to as a two-fluid model with spin-angular momentum conservation. More importantly, our two-fluid model differs from conventional models in terms of the physical picture. The conventional two-fluid model was derived by combining the nonlinear $\rm Schr\ddot{o}dinger$ equation for bosons with the thermodynamic Gibbs-Duhem equation. Conversely, our model considers the bulk of liquid helium-4 at low temperatures as an inviscid fluid following continuum mechanics, which contains residual viscosity at finite temperatures. In summary, the proposed model is based on a two-phase flow. Despite the fully classical mechanical picture, the resulting system equations were consistent with an ordinary model (see Eqs. (1) and (2)). In particular, we consider the bulk liquid helium-4 to be an inviscid fluid, assuming that the viscous fluid component remains at finite temperatures. As the temperature decreased, the amount of the viscous fluid component decreased, ultimately becoming a fully inviscid fluid at absolute zero. From the Lagrangian perspective, our two-fluid model is a mixture of fluid particles from inviscid and viscous fluid components. Both fluid particles are classical fluid particles; they are virtual particles. Thus, our model is based on the mixing of two classical fluids and is similar to a two-phase flow model. However, as described below, we allow for a certain “nature of particles” of both components as a correction. In this context, the physical picture in which the viscous fluid component remains bead-like in an inviscid fluid is a more accurate representation of the system than a two-phase flow model. This is also close to the image of a “nanofluid,” or nanoparticle dispersion system. Details of this technique are described in Refs. [6, 7, 8]. This will be outlined later in this paper. In our previous simulations of rotating two-dimensional (2D) liquid helium-4, we observed the formation of multiple spinning vortices that occurred independently of the overall fluid motion. After incorporating the vortex dynamics model into our model, we found that multiple spinning vortices formed a lattice with a regular arrangement. The number of vortices generated is determined by the strength of the rotational angular velocity. For appropriate values, the number of vortices agrees well with the theoretical solution of the number of vortices calculated from Feynman’s law. It is also noteworthy that the spinning of each vortex moves independently without affecting the behavior of the entire fluid. In other words, the vortices do not dissipate. In a quantum vortex, the vortices do not dissipate because of the circulation quantization. This important property can also be reproduced. Consequently, the vortex lattice phenomenon in the rotating 2D liquid helium-4 was simulated by solving the two-fluid model with spin angular momentum conservation.
@@ 2.1 The two-fluid model with spin angular momentum conservation
Our two-fluid model differs from the ordinary two-fluid model in its derivation and interpretation of the equation. Here, the ordinary two-fluid model is derived from the following microscopic relations. More specifically, the Navier-Stokes equation for an incompressible fluid with a temperature gradient term was derived by solving the Gross-Pitaevskii equation, which is a nonlinear Schrödinger equation, and the Gibbs-Duhem thermodynamic relation in conjunction. This equation describes the dynamics of the superfluid component. In short, the usual two-fluid model is obtained by making “macro-corrections” to the microscopic relational equations. The governing equation for the normal flow component was obtained by subtracting the equation for the superfluid component from the equation of motion for an incompressible fluid. In a domain where quantum mechanics is dominant, the superfluid and normal flow components exist independently, and the two components together are considered to conserve mass and momentum (although entropy is said to be carried only by the normal flow component). This microscopic picture was proposed by Tiza and Landau [15, 16] and was subsequently verified by countercurrent flow experiments. Originally, the two components were considered completely independent; however, experiments have shown that when the heat input is large, the two components exert a mutual frictional force proportional to the velocity difference between them  [17].
@@
By contrast, the concept and derivation of our two-fluid model are quite the opposite. Notably, we obtain a system of two equations of the same form as in the two-fluid model described above by making particle corrections to the fluid equations, which are macroscopic expressions of relations in the domain dominated by continuum mechanics. The details are as follows. First, bulk liquid helium at cryogenic temperatures is considered an inviscid fluid. Subsequently, we assume that the viscous fluid component remains at a finite temperature. The amount of the viscous fluid component decreased as the temperature decreased; at absolute zero, it became a completely inviscid fluid. In other words, our two-fluid model assumes the mixing of the fluid components of inviscid and viscous fluids. Thus, the “fluid particles” in our model are virtual particles in classical fluid mechanics. Liquid helium, on the other hand, is usually considered an incompressible fluid. However, helium atoms are known to cause Bose-Einstein condensations (BECs) when cooled to the critical temperature (interestingly, although liquid helium is commonly described as a multiparticle system of Bose particles, only 13 % of it actually causes BEC even at absolute zero [18]). In any case, the condensate components were no longer fluid. These state changes to and from the condensates are described using quantum statistical mechanics. In other words, from a microscopic viewpoint, the composition ratio of the condensate component to the normal fluid component fluctuates around the average value. Therefore, it can be assumed that the fluid volume fluctuated around the mean value of the equilibrium state. From a macroscopic perspective, the bulk of liquid helium-4 was compressed by a few percent to approximately 13% in the low-temperature regime, and the volume changed around the average. Accordingly, this physical picture is compatible with the fluid computational model of the explicit SPH method. The explicit SPH method uses a finite-particle approximation of the incompressible Navier-Stokes equations. However, instead of implicitly solving the Poisson equation for pressure, as is typically done for incompressible fluids, this method solves the equation of state of a fluid, for example, the Tait equation of state [19]. This approach essentially reproduces density fluctuations [20, 21, 22]. An important point is that we assumed weak compressibility for liquid helium. The ability to solve the Navier-Stokes equations for an incompressible fluid while reproducing its weakly compressible properties is a feature of explicit SPH. In summary, in our two-fluid model, liquid helium is considered an inviscid fluid, and it is assumed that the viscous fluid component remains a function of temperature. In other words, it is a mixture of inviscid and viscous fluids. In addition, weak compressibility was observed. Weak compressibility was reproduced using the explicit SPH method.
Figure 1: Schematic of our methodology: a classical mechanical interpretation of bulk liquid helium-4
@@
Figure 1 presents a schematic of the proposed method. Although the two components may exhibit particle properties as described above, in principle, they are “fluid components,” and therefore, an interface between the two components is considered to exist under certain conditions. Even if the temperature of the fluid is constant, the density difference causes a temperature gradient between the two fluid particles near the interface, which acts as interfacial tension. In summary, an interfacial tension proportional to the temperature gradient is generated in the opposite direction between the two fluid components, as shown in the center of Fig.  1. In summary, the inviscid component follows the Euler equation with a temperature gradient term, whereas the viscous component follows the incompressible Navier–Stokes equation with a temperature gradient term in the opposite direction. At this stage, only the temperature gradient force contributes to the interfacial tension between the two components. Temperature is a macroscopic state quantity. In other words, only the macroscopic physical properties contribute to interfacial tension. However, as the temperature approaches absolute zero, more local interactions between the two components are expected to contribute to interfacial tension. In other words, local properties have emerged as multiparticle systems. Therefore, in addition to the macroscopic temperature gradient force, the local interaction force $\bm{\mathbf{F}}_{sn}$ between two fluid particles is explicitly described as part of the interfacial tension. In classical fluid mechanics, each phase is subjected to an action proportional to the local velocity difference between the different phases at the interface of a two-phase flow. In the Lagrangian picture, the action force between two different fluid particles is proportional to the local velocity difference between the two components. More specifically, as shown on the left-hand side of Fig. 1, the $j$th fluid particle is subjected to a force proportional to the velocity difference between its own velocity and the average of the velocities of different types of fluid particles, and its reaction force is reflected in the different types of fluid particles. Because we use the explicit SPH method described above, the interface model of the two-layer flow of the explicit SPH method can be used as is. We examined the conditions under which the interface between the two phases should be considered. First, the “interface” between the two phases is a fundamental concept in hydrodynamics. Therefore, the interface effect becomes non-negligible when the two components exhibit hydrodynamic rather than many-body properties. Liquid helium-4 is expected to behave as a uniform inviscid when no external force is applied. However, the assumption of a single-phase flow of an inviscid fluid is broken by a strong external shock or heat, and an interface effect between the two components appears. As this situation simultaneously reveals the multi-particle nature of the fluid, the interfacial effect operates only in the direction of the applied impact force, and the force does not propagate in other directions as in a continuum. In the horizontal rotation problem, liquid helium-4 was forced to rotate horizontally; therefore, fluidity with weak compressibility may manifest strongly in the horizontal direction. Accordingly, the driving force due to the temperature gradient force was assumed to act in the horizontal direction. Interface effects in this direction were not considered in our model because no external forces acted in the vertical direction.
@@
The above can be summarized as follows. Our two-fluid model assumes mixing of inviscid and viscous fluid components. The inviscid component follows the Euler equation with a temperature gradient term, whereas the viscous component follows the incompressible Navier-Stokes equation with a temperature gradient term in the opposite direction. The interface effect between the two components owing to a temperature gradient acts in the induced direction only when classical fluidity is induced between the two components by excitation or external forces. We consider the local interaction force $\bm{\mathbf{F}}_{sn}$ acting between two different types of fluid particles that is proportional to the velocity difference between the different types of fluid particles. In addition, the viscous term was reformulated to conserve the angular momentum of the fluid particles around their axes, that is, the spin angular momentum, as previously described. The resulting system of equations is as follows [6, 7, 8]
@@
$\rho_{n}\frac{{\rm D}\bm{\mathbf{v}}_{n}}{{\rm D}t}$ $=$ $-\frac{\rho_{n}}{\rho}\nabla P-\rho_{s}\sigma\nabla T+(\bar{\eta}+\bar{\eta_{r}})\nabla^{2}\bm{\mathbf{v}}_{n}+\biggl{(}\frac{\bar{\eta}}{3}+\bar{\xi}-\bar{\eta_{r}}\biggr{)}\nabla\nabla\cdot\bm{\mathbf{v}}_{n}+2\bar{\eta_{r}}\nabla\times\bm{\mathbf{\omega}}+\bm{\mathbf{F}}_{sn}.$ (2)

### v2

@@ Abstract
This paper reports a three-dimensional (3D) simulation of a rotating liquid helium-4, using a two-fluid model with spin-angular momentum conservation. Our model was derived from the particle approximation of an inviscid fluid with residual viscosity. Despite the fully classical mechanical picture, the resulting system equations were consistent with those of the conventional two-fluid model. We consider bulk liquid helium-4 to be an inviscid fluid, assuming that the viscous fluid component remains at finite temperatures. As the temperature decreased, the amount of the viscous fluid component decreased, ultimately becoming a fully inviscid fluid at absolute zero. Weak compressibility is assumed to express the volume change because some helium atoms do not render fluid owing to Bose-Einstein condensations (BECs) or change states because of local thermal excitation. One can solve the governing equations for an incompressible fluid using explicit SPH (smoothed-particle hydrodynamics), simultaneously reproducing density fluctuations and describing the fluid in a many-particle system. We assume the following fluid-particle duality: a hydrodynamic interfacial tension between the inviscid and viscous components or a local interaction force between two types of fluid particles. The former can be induced in the horizontal direction when non-negligible non-uniformity of the particles occurs during forced two-dimensional rotation, and the latter is non-negligible when the former is negligible. We performed a large-scale simulation of 3D liquid helium forced to rotate horizontally using 32 number of graphics processing units (GPUs). Compared with the low-resolution calculation using 2.4 million particles, the high-resolution calculation using 19.6 million particles showed spinning vortices close to those of the theoretical solution. We obtained a promising venue to establish a practical simulation method for bulk liquid helium-4.
@@ 1 Introduction
Considering this hypothesis, the author recently developed a two-fluid model with spin-angular momentum conservation [6, 7, 8]. The following overview provides a general introduction to this subject. The Navier-Stokes equation with conservation of rotational angular momentum around the axes of fluid particles was originally derived from the constitutive law of continuum mechanics by D. W. Condiff in 1964 [9]. Half a century later, K. Müller [10] developed a particle approximation of the Navier-Stokes equation with spin-angular momentum conservation using smoothed-particle hydrodynamics (SPH) [11, 12] for mesoscale thermal flow. Considering these studies, the author adopted the Navier-Stokes equations with spin-angular momentum conservation for normal fluids and employed SPH to discretize the equation systems using particles. In other words, we describe liquid helium-4 as a system of numerous particle interactions composed of two distinct types of classical fluid particles that maintain their rotational angular velocity around their axes under the assumption of rigid body rotations; this is referred to as a two-fluid model with spin-angular momentum conservation. More importantly, our two-fluid model differs from conventional models in terms of the physical picture. The conventional two-fluid model was derived by combining the nonlinear $\rm Schr\ddot{o}dinger$ equation for bosons with the thermodynamic Gibbs–Duhem equation. Conversely, our model considers the bulk of liquid helium-4 at low temperatures as an inviscid fluid following continuum mechanics, which contains residual viscosity at finite temperatures. In summary, the proposed model is based on a two-phase flow. Despite the fully classical mechanical picture, the resulting system equations were consistent with an ordinary model (see Eqs. (1) and (2)). In particular, we consider the bulk liquid helium-4 to be an inviscid fluid, assuming that the viscous fluid component remains at finite temperatures. As the temperature decreased, the amount of the viscous fluid component decreased, ultimately becoming a fully inviscid fluid at absolute zero. From the Lagrangian perspective, our two-fluid model is a mixture of fluid particles from inviscid and viscous fluid components. Both fluid particles are classical fluid particles; they are virtual particles. Thus, our model is based on the mixing of two classical fluids and is similar to a two-phase flow model. However, as described below, we allow for a certain “nature of particles” of both components as a correction. In this context, the physical picture in which the viscous fluid component remains bead-like in an inviscid fluid is a more accurate representation of the system than a two-phase flow model. This is also close to the image of a “nanofluid,” or nanoparticle dispersion system, where the nanoparticles maintain fluidity. Details of this technique are described in Refs. [6, 7, 8]. This will be outlined later in this paper. In our previous simulations of rotating two-dimensional (2D) liquid helium-4, we observed the formation of multiple spinning vortices that occurred independently of the overall fluid motion. After incorporating the vortex dynamics model into our model, we found that multiple spinning vortices formed a lattice with a regular arrangement. The number of vortices generated is determined by the strength of the rotational angular velocity. For appropriate values, the number of vortices agrees well with the theoretical solution of the number of vortices calculated from Feynman’s law. It is also noteworthy that the spinning of each vortex moves independently without affecting the behavior of the entire fluid. In other words, the vortices do not dissipate. In a quantum vortex, the vortices do not dissipate because of the circulation quantization. This important property can also be reproduced. Consequently, the vortex lattice phenomenon in the rotating 2D liquid helium-4 was simulated by solving the two-fluid model with spin angular momentum conservation.
@@ 2.1 The two-fluid model with spin angular momentum conservation
Our two-fluid model differs from the ordinary two-fluid model in its derivation and interpretation of the equation. Here, the ordinary two-fluid model is derived from the following microscopic relations. More specifically, the Navier-Stokes equation for an inviscid fluid with a temperature gradient term was derived by solving the Gross–Pitaevskii equation, which is a nonlinear $\rm Schr\ddot{o}dinger$ equation, and the Gibbs–Duhem thermodynamic relation in conjunction. This equation describes the dynamics of the superfluid component. In short, the usual two-fluid model is obtained by making “macro-corrections” to the microscopic relational equations. The governing equation for the normal flow component was obtained by subtracting the equation for the superfluid component from the momentum balance equation of the total fluid [15]. Figure 1(II) schematically shows the relationship among the Gross–Pitaevskii equation, Gibbs–Duhem thermodynamic relation, and the ordinary two-fluid model. In a domain where quantum mechanics is dominant, the superfluid and normal flow components exist independently, and the two components together are considered to conserve mass and momentum (although entropy is said to be carried only by the normal flow component). This microscopic picture was proposed by Tiza and Landau [16, 17] and was subsequently verified by countercurrent flow experiments. Originally, the two components were considered completely independent; however, experiments have shown that when the heat input is large, the two components exert a mutual frictional force proportional to the velocity difference between them  [18].
@@
By contrast, the concept and derivation of our two-fluid model are quite the opposite. Notably, we obtain a system of two equations of the same form as in the two-fluid model described above by making particle corrections to the fluid equations, which are macroscopic expressions of relations in the domain dominated by continuum mechanics. Figure 1(I) presents a schematic of our methodology. The details are as follows. First, bulk liquid helium at cryogenic temperatures is considered an inviscid fluid. Subsequently, we assume that the viscous fluid component remains at a finite temperature, as illustrated in Fig. 1(a). The amount of the viscous fluid component decreased as the temperature decreased; at absolute zero, it became a completely inviscid fluid. In other words, our two-fluid model assumes the mixing of the fluid components of inviscid and viscous fluids. Thus, the “fluid particles” in our model are virtual particles in classical fluid mechanics. Liquid helium, on the other hand, is usually considered an incompressible fluid. However, helium atoms are known to cause Bose-Einstein condensations (BECs) when cooled to the critical temperature (interestingly, although liquid helium is commonly described as a multiparticle system of Bose particles, only 13% of it actually causes BEC even at absolute zero [19]). In any case, the condensate components were no longer fluid. These state changes to and from the condensates are described using quantum statistical mechanics. In other words, from a microscopic viewpoint, the composition ratio of the condensate component to the normal fluid component fluctuates around the average value. Therefore, it can be assumed that the fluid volume fluctuated around the mean value of the equilibrium state. From a macroscopic perspective, the bulk of liquid helium-4 was compressed by a few percent to approximately 13% in the low-temperature regime, and the volume changed around the average. Accordingly, this physical picture is compatible with the fluid computational model of the explicit SPH method. The explicit SPH method uses a finite-particle approximation of the incompressible Navier-Stokes equations. However, instead of implicitly solving the Poisson equation for pressure, as is typically done for incompressible fluids, this method solves the equation of state of a fluid, for example, the Tait equation of state [20]. This approach essentially reproduces density fluctuations [21, 22, 23]. An important point is that we assumed weak compressibility for liquid helium. The ability to solve the Navier-Stokes equations for an incompressible fluid while reproducing its weakly compressible properties is a feature of explicit SPH. In summary, in our two-fluid model, liquid helium is considered an inviscid fluid, and it is assumed that the viscous fluid component remains a function of temperature. In other words, it is a mixture of inviscid and viscous fluids. In addition, weak compressibility was observed. Weak compressibility was reproduced using the explicit SPH method.
Figure 1: Schematic of our methodology: (I) a classical mechanical interpretation of bulk liquid helium-4, and (II) the derivation of the conventional two-fluid model in quantum hydrodynamics for comparison. Each part presents schematics of (a) the mixture of inviscid and viscid fluids, (b) the interfacial effect due to the temperature gradient induced under certain conditions, (c) the local interaction forces between different types of fluid particles, (d) the spin-angular momentum conservation, and (e) the use of the explicit SPH to realize the density fluctuation (weak compressibility). Next, (f) visually illustrates the derivation of the original two-fluid model from the Gross–Pitaevskii equation and the Gibbs–Duhem thermodynamic relation, (g) describes the introduction of mutual frictional forces. (h) is the Euler equation for an inviscid fluid, (i) is the Navier-Stokes equation for a viscous fluid, and (j) is a reformulated version of (i) introducing spin-angular momentum conservation; (h)-(j) are derived from the classical mechanical assumption. In contrast, (k) and (l) represent the ordinary two-fluid model in quantum hydrodynamics.
@@
Although the two components may exhibit particle properties as described above, in principle, they are “fluid components,” and therefore, an interface between the two components is considered to exist under certain conditions. Even if the temperature of the fluid is constant, the density difference causes a temperature gradient between the two fluid particles near the interface, which acts as interfacial tension. In summary, an interfacial tension proportional to the temperature gradient is generated in the opposite direction between the two fluid components, as shown in Fig. 1(b). In summary, the inviscid component follows the Euler equation with a temperature gradient term, whereas the viscous component follows the incompressible Navier–Stokes equation with a temperature gradient term in the opposite direction (Figs. 1(a) and (b)). At this stage, only the temperature gradient force contributes to the interfacial tension between the two components. Temperature is a macroscopic state quantity. In other words, only the macroscopic physical properties contribute to interfacial tension. However, as the temperature approaches absolute zero, more local interactions between the two components are expected to contribute to interfacial tension. In other words, local properties have emerged as multiparticle systems. Therefore, in addition to the macroscopic temperature gradient force, the local interaction force $\bm{\mathbf{F}}_{sn}$ between two fluid particles is explicitly described as part of the interfacial tension. In classical fluid mechanics, each phase is subjected to an action proportional to the local velocity difference between the different phases at the interface of a two-phase flow. In the Lagrangian picture, the action force between two different fluid particles is proportional to the local velocity difference between the two components. More specifically, as shown in Fig. 1(c), the $i$th fluid particle is subjected to a force proportional to the velocity difference between its own velocity and the average of the velocities of different types of fluid particles, and its reaction force is reflected in the different types of fluid particles. Because we use the explicit SPH method described above, the interface model of the two-layer flow of the explicit SPH method can be used as is. We examined the conditions under which the interface between the two phases should be considered. First, the “interface” between the two phases is a fundamental concept in hydrodynamics. Therefore, the interface effect becomes non-negligible when the two components exhibit hydrodynamic rather than many-body properties. Liquid helium-4 is expected to behave as a uniform inviscid fluid when no external force is applied. However, the assumption of a single-phase flow of an inviscid fluid is broken by a strong external shock or heat, and an interface effect between the two components appears. As this situation simultaneously reveals the multi-particle nature of the fluid, the interfacial effect operates only in the direction of the applied impact force, and the force does not propagate in other directions as in a continuum. In the horizontal rotation problem, liquid helium-4 was forced to rotate horizontally; therefore, fluidity with weak compressibility may manifest strongly in the horizontal direction. Accordingly, the driving force due to the temperature gradient force was assumed to act in the horizontal direction. Interface effects in the vertical direction were not considered in our model because no external forces acted in this direction.
@@
The above can be summarized as follows. Our two-fluid model assumes mixing of inviscid and viscous fluid components. The inviscid component follows the Euler equation with a temperature gradient term, whereas the viscous component follows the incompressible Navier-Stokes equation with a temperature gradient term in the opposite direction. The interface effect between the two components owing to a temperature gradient acts in the induced direction only when classical fluidity is induced between the two components by excitation or external forces. We consider the local interaction force $\bm{\mathbf{F}}_{sn}$ acting between two different types of fluid particles that is proportional to the velocity difference between the different types of fluid particles. In addition, as shown in Fig. 1(d), the viscous term was reformulated to conserve the angular momentum of the fluid particles around their axes, that is, the spin angular momentum. This is explained further later. We admit weak compressibility under specific conditions, and therefore, adopt an explicit SPH as in Fig. 1(e). The resulting system of equations is as follows [6, 7, 8]
@@
$\rho_{n}\frac{{\rm D}\bm{\mathbf{v}}_{n}}{{\rm D}t}$ $=$ $-\frac{\rho_{n}}{\rho}\nabla P-\rho_{s}\sigma\nabla T+(\bar{\eta}+\bar{\eta_{r}})\nabla^{2}\bm{\mathbf{v}}_{n}+\biggl{(}\frac{\bar{\eta}}{3}+\bar{\xi}-\bar{\eta_{r}}\biggr{)}\nabla\nabla\cdot\bm{\mathbf{v}}_{n}+2\bar{\eta_{r}}\nabla\times\bm{\mathbf{\omega}}_{0}+\bm{\mathbf{F}}_{sn}.$ (2)
As shown in Eqs. (1) and (2), the mathematical expressions for the five assumptions shown in Figs. 1(a)–(e) can be described in a form similar to that of the conventional two-fluid model derived from the $\rm Schr\ddot{o}dinger$ equation. Let us explain this further. Figure 1(I) illustrates the assumptions (a)–(e) that derive our two-fluid model. Equations (h) and (j) in Fig. 1 correspond to Eqs. (1) and (2), respectively. The underlined subscripts in each term of the equations indicate the assumptions underlying the introduction of these terms. For example, Eq. (h) is the Euler equation for an inviscid fluid with a temperature gradient and is based on assumptions (a)–(c). Equation (i) can be derived based on assumptions (a)–(c), and it can also be obtained by subtracting Eq. (h) from the momentum balance equation for the total fluid [15]. In addition, we reformulate the viscosity term in Eq. (i) to conserve the rotational angular momentum of each constituent particle, as described below. For comparison, the ordinary two-fluid model, which was derived from the Gross–Pitaevskii and Gibbs–Duhem thermodynamic equations, is shown in (k) and (l). As previously mentioned, Eq. (l) can be obtained by subtracting Eq. (k) from the momentum balance equation for the total fluid [15]. In other words, the only difference between the derivation of the two-fluid model in (I) classical and (II) microscopic depictions is whether the existence of the inviscid fluid is a precondition or whether the inviscid flow is also derived from the microscopic equations and thermodynamic relations. However, the importance of these differences in interpretation has a significant impact on subsequent modeling; this is discussed in section 2.2. To summarize, it can be seen from Fig. 1 that the governing equations for cryogenic liquid helium are expressed in similar forms both from the classical fluid or quantum mechanical perspectives.